OpenStudy (anonymous):

Solve, using linear combination. 3x + y = 4 2x + y = 5

OpenStudy (anonymous):

3x + y =4 2x +y = 5 then y=5-2x combine 2 expression 3x +(5-2x) = 4 then x= -1 y = 5-2*(-1) y=7